OpenStudy (anonymous):

11%of what number is 0.7

OpenStudy (anonymous):

Make it an algebraic equation! - Of means Multiplication - % means Divided by 100 - Let the unknown number be x Therefore, (11/100)*x=0.7

OpenStudy (anonymous):

is over of equals % over 100. 0.7 over x equals 11 over 100. cross multiply.

OpenStudy (anonymous):

x=0.7/0.11 x=6.363636363636364